Using laws of exponents, simplify and write the answer in exponential form:  6^15 ÷ 6 ^10

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

By the rule of dividing the powers with same base = a^m ÷a^n= a^m - n
Then,
= (6)^15 - 10
= 6^5

Thanks